Solve the stem of linear equations. What is the x and y answer to this problem?
2x + y = 4
x - y = 2

Answer by mananth(16946)

2 x + 1 y = 4 .............1
1 x -1 y = 2 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
2 x 1 y = 4
1 x -1 y = 2
Add the two equations
3 x = 6
/ 3
x = 2
plug value of x in (1)
2 x + 1 y = 4
4 + 1 y = 4
1 y = 4 -4
1 y = 0
y = 0